# Exponential actions defined by vector configurations, Gale duality, and moment-angle manifolds

### v1

@@ 11. Gale duality for rational configurations
A simple diagram chase (see [13, Lemma 1.2.5]) shows that $M\xrightarrow{{\widehat{A}}^{*}}\mathbb{Z}^{m}\to\mathbb{Z}^{I}$ is surjective if and only if $\mathbb{Z}^{\widehat{I}}\to\mathbb{Z}^{m}\xrightarrow{\widehat{\varGamma}}L^{*}$ is surjective. The latter is equivalent to the condition that $\widehat{\Gamma}_{\widehat{I}}$ spans $L^{*}$. ∎
@@
A cone in a vector space $U^{*}$ with a full-rank lattice $M^{*}$ is called rational (with respect to $M^{*}$) if it is generated by a set of vectors in $M^{*}$. A rational cone in nonsingular if it is generated by part of a basis of $M^{*}$. A nonsingular cone is necessarily simplicial. A fan is rational (respectively, nonsingular) if it is consists of rational (respectively, nonsigular) cones.
@@
Let $\Gamma$ and $\mathrm{A}$ be a pair of Gale dual vector configurations, $Q\subset V$ a rational subspace, and $\mathcal{K}$ a simplicial complex on $[m]$. Assume that $\{\mathcal{K},\mathrm{A}\}$ defines a simplicial fan $\Sigma$ in $W^{*}$. Then
@@
    $\{\mathcal{K},\widehat{\mathrm{A}}\}$ defines a rational simplicial fan $\widehat{\Sigma}$ in $U^{*}$.
@@
Since $\{\mathcal{K},\mathrm{A}\}$ defines a fan, we have $(\mathop{\mathrm{relint}}\mathop{\mathrm{cone}}\mathrm{A}_{I})\cap(\mathop{\mathrm{relint}}\mathop{\mathrm{cone}}\mathrm{A}_{J})=\varnothing$ for any $I,J\in\mathcal{K}$, $I\neq J$. By Proposition 4.1, the surjective map $U^{*}\to W^{*}$ takes $\mathop{\mathrm{relint}}\mathop{\mathrm{cone}}\widehat{\mathrm{A}}_{I}$ to $\mathop{\mathrm{relint}}\mathop{\mathrm{cone}}\mathrm{A}_{I}$. Hence, $(\mathop{\mathrm{relint}}\mathop{\mathrm{cone}}\widehat{\mathrm{A}}_{I})\cap(\mathop{\mathrm{relint}}\mathop{\mathrm{cone}}\widehat{\mathrm{A}}_{J})=\varnothing$ for any $I,J\in\mathcal{K}$, $I\neq J$. It follows that the data $\{\mathcal{K},\widehat{\mathrm{A}}\}$ also defines a fan, proving (1). Statement (2) is clear, as all the fans involved are simplicial. ∎
@@
Assume that $\{\mathcal{K},\widehat{\mathrm{A}}\}$ defines a fan $\widehat{\Sigma}$ in $U^{*}$. Then $\widehat{\Sigma}$ is nonsingular if and only if $L_{I}=L$ for any $I\in\mathcal{K}$.
@@
Assume that $L_{I}=L$. Then $\widehat{\Gamma}_{\widehat{I}}=\{\widehat{\gamma}_{k}\colon k\notin I\}$ spans $Q^{*}$. Indeed, otherwise there is ${\textbf{{q}}}\in Q$ such that $\langle\widehat{\gamma}_{k},{\textbf{{q}}}\rangle=0$ for $k\notin I$, so that $L_{I}$ contains the subspace $\mathbb{R}\langle q\rangle$. On the other hand, $L$ is a lattice and cannot contain a subspace. Now $\mathbb{R}\langle\widehat{\Gamma}_{\widehat{I}}\rangle=Q^{*}$ implies that $\mathbb{Z}\langle\widehat{\Gamma}_{\widehat{I}}\rangle$ is a lattice in $Q^{*}$ whose dual lattice is $L_{I}$. We obtain $\mathbb{Z}\langle\widehat{\Gamma}_{\widehat{I}}\rangle=L_{I}^{*}=L^{*}$ for any $I\in\mathcal{K}$. Then $\widehat{\Sigma}$ is a nonsingular fan by Proposition 11.1.
Now assume that $\widehat{\Sigma}$ is a nonsingular fan. Then $\mathbb{Z}\langle\widehat{\Gamma}_{\widehat{I}}\rangle=L^{*}=\mathbb{Z}\langle\widehat{\Gamma}\rangle$ for any $I\in\mathcal{K}$ by Proposition 11.1. It follows that each $\widehat{\gamma}_{i}$, $i=1,\ldots,m$, is an integral linear combination of $\{\widehat{\gamma}_{k}\colon k\notin I\}$, so that $L_{I}=L$. ∎
@@ 12. Partial quotients and torus-exponential actions
A partial quotient of a moment-angle manifold $\mathcal{Z}_{\mathcal{K}}$ is the quotient $\mathcal{Z}_{\mathcal{K}}/H$ by a freely acting subtorus $H\subset T^{m}$. When $\{\mathcal{K},\mathrm{A}\}$ is a fan data and the subtorus $H$ is defined by a rational subspace $Q\subset V$ of even codimension, the partial quotient $\mathcal{Z}_{\mathcal{K}}/H$ admits a complex structure in which the quotient torus $T^{m}/H$ acts by holomorphic transformations. Such a holomorphic partial quotient is the quotient of the open subset $U_{\mathbb{C}}(\mathcal{K})$ by a torus-exponential action of the product of an algebraic torus and a complex space. In addition to complex moment-angle manifolds, this construction includes LVMB manifolds and complete nonsingular toric varieties as special cases.
@@ 12.1. Torus-exponential actions and their quotients
Given a vector configuration $\Gamma=\{\gamma_{1},\ldots,\gamma_{m}\}$ in $V$, the map $\varGamma^{*}|_{L}\otimes_{\mathbb{Z}}\mathbb{C}^{\times}$ embeds $\mathbb{C}^{\times}_{L}$ as a closed (algebraic) subgroup in (×$(\mathbb{C}^{\times})^{m}$:
@@
An action of a Lie group is almost free if all stabilizers are discrete subgroups.
@@
Let $\Gamma$ and $\mathrm{A}$ be a pair of Gale dual vector configurations, $Q\subset V$ a rational subspace, and $\mathcal{K}$ a simplicial complex on $[m]$. The torus-exponential action $(\mathbb{C}^{\times}_{L}\times_{Q}V)\times U_{\mathbb{C}}(\mathcal{K})\to U_{\mathbb{C}}(\mathcal{K})$ is proper if and only if $\{\mathcal{K},\mathrm{A}\}$ is a simplicial fan data, in which case the action is almost free with finite stabilizers.
@@
Since $T_{Q}$ is a compact group, the action of $\mathbb{C}^{\times}_{L}\times_{Q}V\cong T_{Q}\times V$ is proper if and only if the action of $V$ is proper. Hence, the assertion about the properness follows from Theorem 5.1. It also gives that the action of $V$ is free. For the action of $T_{L}$, the stabilizer of ${\textbf{{z}}}\in U_{\mathbb{C}}(\mathcal{K})$ is $2\pi(L_{I}/L)$. Since $I\in\mathcal{K}$, the set $\{\gamma_{k}\colon k\notin I\}$ spans $V^{*}$, so $L_{I}$ is a full-rank lattice in $Q$. It follows that $L_{I}/L$ is a finite group. ∎
@@
Let $\Gamma$ and $\mathrm{A}$ be a pair of Gale dual vector configurations, $Q\subset V$ a rational subspace, and $\mathcal{K}$ a simplicial complex on $[m]$. If $\{\mathcal{K},\mathrm{A}\}$ is a simplicial fan data, then the action of $\mathbb{C}^{\times}_{L}$ on $U_{\mathbb{C}}(\mathcal{K})$ is proper.
@@
is proper. This is the exponential action defined by $\widehat{\Gamma}=\{\gamma_{1},\ldots,\gamma_{m}\}$ viewed as a vector configuration in the quotient space $Q^{*}$ of $V^{*}$. Let $\widehat{\mathrm{A}}$ be the Gale dual configuration. Then $\{\mathcal{K},\widehat{\mathrm{A}}\}$ is a fan data by Proposition 11.3. It follows that the action is proper, by the criterion of Theorem 4.4. ∎
The quotient $U_{\mathbb{C}}(\mathcal{K})/\mathbb{C}^{\times}_{L}$ by a proper algebraic action of $\mathbb{C}^{\times}_{L}$ is the toric variety $X_{\widehat{\Sigma}}$ corresponding to the rational simplicial fan $\widehat{\Sigma}$ defined by $\{\mathcal{K},\widehat{\mathrm{A}}\}$ [15, §5.1].
@@ 12.2. Holomorphic torus-exponential actions
Let $\Gamma$ and $\mathrm{A}$ be a pair of Gale dual vector configurations, $Q\subset V$ a rational subspace, and $\mathcal{K}$ a simplicial complex on $[m]$. Suppose that
@@
    the data $\{\mathcal{K},\mathrm{A}\}$ defines a simplicial fan $\Sigma$;
@@
    the fan $\widehat{\Sigma}$ defined by $\{\mathcal{K},\widehat{\mathrm{A}}\}$ is nonsingular;
@@
Condition (a) implies that the action of $\mathbb{C}^{\times}_{L}\times\widetilde{V/Q}$ on $U_{\mathbb{C}}(\mathcal{K})$ is proper. Condition (b) together with Propositions 11.4 and 12.1 implies that the action of $T_{L}$ on $U_{\mathbb{C}}(\mathcal{K})$ is free, hence, the action of $\mathbb{C}^{\times}_{L}\times\widetilde{V/Q}$ is also free. Finally, the homeomorphism in the bottom right conner is proved in the same way as Theorem 10.6 (which corresponds to the case $Q=\{\mathbf{0}\}$). ∎
@@
If the rational fan $\widehat{\Sigma}$ is nonsingular, then the actions of $\mathbb{C}^{\times}_{L}$ and $\widetilde{T}_{L}$ are free, and $\mathcal{Z}_{\mathcal{K}}\to\mathcal{Z}_{\mathcal{K}}/\widetilde{T}_{L}$ is a holomorphic torus principal bundle over the holomorphic partial quotient. More generally, $\mathcal{Z}_{\mathcal{K}}\to\mathcal{Z}_{\mathcal{K}}/\widetilde{T}_{L}$ is a holomorphic Seifert principal bundle [30, §5.1] over the orbifold $\mathcal{Z}_{\mathcal{K}}/\widetilde{T}_{L}\cong X_{\widehat{\Sigma}}/(\widetilde{V/Q})$, where the latter the quotient of the toric orbifold $X_{\widehat{\Sigma}}$ by a free proper action of $\widetilde{V/Q}$.
@@ 12.3. Maximal torus actions
An action of a torus $T$ on a connected smooth manifold $M$ is maximal if there is a point $x\in M$ such that $\dim T+\dim T_{x}=\dim M$, where $T_{x}$ is the stabiliser subgroup of $x$. Examples of maximal torus actions are the action of a half-dimensional compact torus on a smooth toric variety, the free action of $T^{m}$ on itself and the action of $T^{m}$ on a moment-angle manifold $\mathcal{Z}_{\mathcal{K}}$. These are particular cases of the action of $T^{m}/T_{L}$ on a partial quotient $\mathcal{Z}_{\mathcal{K}}/T_{L}$:
@@
Compact complex manifolds $M$ with maximal torus actions by holomorphic transformations are classified by Ishida in [21]. The main result [21, Theorem 1.5] implies that every such $M$ is equivariantly biholomorphic to a holomorphic partial quotient $\mathcal{Z}_{\mathcal{K}}/T_{L}$. (For the latter formulation, see [36, Theorem 4.8].) It also implies that any complex manifold $M$ with a maximal torus action is the base of a holomorphic torus principal bundle with total space a complex moment-angle manifold.

### v2

@@ 11. Gale duality for rational configurations
A simple diagram chase (see [14, Lemma 1.2.5]) shows that $M\xrightarrow{{\widehat{A}}^{*}}\mathbb{Z}^{m}\to\mathbb{Z}^{I}$ is surjective if and only if $\mathbb{Z}^{\widehat{I}}\to\mathbb{Z}^{m}\xrightarrow{\widehat{\varGamma}}L^{*}$ is surjective. The latter is equivalent to the condition that $\widehat{\Gamma}_{\widehat{I}}$ spans $L^{*}$. ∎
@@
A cone in a vector space $U^{*}$ with a full-rank lattice $M^{*}$ is called rational (with respect to $M^{*}$) if it is generated by a set of vectors in $M^{*}$. A rational cone in nonsingular if it is generated by part of a basis of $M^{*}$. A nonsingular cone is necessarily simplicial. A fan is rational (respectively, nonsingular) if it is consists of rational (respectively, nonsingular) cones.
A configuration $\mathrm{A}=\{{\textbf{{a}}}_{1},\ldots,{\textbf{{a}}}_{m}\}$ of vectors in $U^{*}$ is rational if each ${\textbf{{a}}}_{i}$ belongs to the lattice $M^{*}$. A triangulated configuration $\{\mathcal{K},\mathrm{A}\}$ is nonsingular if $\{{\textbf{{a}}}_{i}\colon i\in I\}$ is part of a basis of $M^{*}$ for any $I\in\mathcal{K}$. A nonsingular triangulated configuration $\{\mathcal{K},\mathrm{A}\}$ defines a nonsingular fan $\Sigma=\{\mathop{\mathrm{cone}}\mathrm{A}_{I}\colon I\in\mathcal{K}\}$ in which ${\textbf{{a}}}_{i}$ is a primitive generator whenever $\{i\}\in\mathcal{K}$.
@@
Let $\Gamma$ and $\mathrm{A}$ be a pair of Gale dual vector configurations, $Q\subset V$ a rational subspace, $\widehat{\Gamma}$ and $\widehat{\mathrm{A}}$ the corresponding rational configurations in $Q^{*}$ and $U^{*}$, and $\mathcal{K}$ a simplicial complex on $[m]$. Assume that $\{\mathcal{K},\mathrm{A}\}$ is a triangulated configuration defining a simplicial fan $\Sigma$ in $W^{*}$. Then
@@
    $\{\mathcal{K},\widehat{\mathrm{A}}\}$ is a rational triangulated configuration defining a rational simplicial fan $\widehat{\Sigma}$ in $U^{*}$.
@@
Since $\{\mathcal{K},\mathrm{A}\}$ is a triangulated configuration, we have $(\mathop{\mathrm{relint}}\mathop{\mathrm{cone}}\mathrm{A}_{I})\cap(\mathop{\mathrm{relint}}\mathop{\mathrm{cone}}\mathrm{A}_{J})=\varnothing$ for any $I,J\in\mathcal{K}$, $I\neq J$. By Proposition 4.1, the surjective map $U^{*}\to W^{*}$ takes $\mathop{\mathrm{relint}}\mathop{\mathrm{cone}}\widehat{\mathrm{A}}_{I}$ to $\mathop{\mathrm{relint}}\mathop{\mathrm{cone}}\mathrm{A}_{I}$. Hence, $(\mathop{\mathrm{relint}}\mathop{\mathrm{cone}}\widehat{\mathrm{A}}_{I})\cap(\mathop{\mathrm{relint}}\mathop{\mathrm{cone}}\widehat{\mathrm{A}}_{J})=\varnothing$ for any $I,J\in\mathcal{K}$, $I\neq J$. It follows that the data $\{\mathcal{K},\widehat{\mathrm{A}}\}$ is also a triangulated configuration, proving (1). Statement (2) is clear, as all the fans involved are simplicial. ∎
@@
A triangulated configuration $\{\mathcal{K},\widehat{\mathrm{A}}\}$ is nonsingular if and only if $L_{I}=L$ for any $I\in\mathcal{K}$.
@@
Assume that $L_{I}=L$ for any $I\in\mathcal{K}$. Then $\widehat{\Gamma}_{\widehat{I}}=\{\widehat{\gamma}_{k}\colon k\notin I\}$ spans $Q^{*}$. Indeed, otherwise there is ${\textbf{{q}}}\in Q$ such that $\langle\widehat{\gamma}_{k},{\textbf{{q}}}\rangle=0$ for $k\notin I$, so that $L_{I}$ contains the subspace $\mathbb{R}\langle{\textbf{{q}}}\rangle$. On the other hand, $L$ is a lattice and cannot contain a subspace. Now $\mathbb{R}\langle\widehat{\Gamma}_{\widehat{I}}\rangle=Q^{*}$ implies that $\mathbb{Z}\langle\widehat{\Gamma}_{\widehat{I}}\rangle$ is a full-rank lattice in $Q^{*}$ whose dual lattice is $L_{I}$. We obtain $\mathbb{Z}\langle\widehat{\Gamma}_{\widehat{I}}\rangle=L_{I}^{*}=L^{*}$. This implies, by Proposition 11.1, that $\widehat{\mathrm{A}}_{I}$ is part of basis of $M^{*}$ for any $I\in\mathcal{K}$. Hence, the triangulated configuration $\{\mathcal{K},\widehat{\mathrm{A}}\}$ is nonsingular.
Now assume that $\{\mathcal{K},\widehat{\mathrm{A}}\}$ is a nonsingular triangulated configuration. Then $\mathbb{Z}\langle\widehat{\Gamma}_{\widehat{I}}\rangle=L^{*}=\mathbb{Z}\langle\widehat{\Gamma}\rangle$ for any $I\in\mathcal{K}$ by Proposition 11.1. It follows that each $\widehat{\gamma}_{i}$, $i=1,\ldots,m$, is an integral linear combination of $\{\widehat{\gamma}_{k}\colon k\notin I\}$, so that $L_{I}=L$. ∎
@@ 12. Partial quotients and torus-exponential actions
A partial quotient of a moment-angle manifold $\mathcal{Z}_{\mathcal{K}}$ is the quotient $\mathcal{Z}_{\mathcal{K}}/H$ by a freely acting subtorus $H\subset T^{m}$. When $\{\mathcal{K},\mathrm{A}\}$ is triangulated configuration and the subtorus $H$ is defined by a rational subspace $Q\subset V$ of even codimension, the partial quotient $\mathcal{Z}_{\mathcal{K}}/H$ admits a complex structure in which the quotient torus $T^{m}/H$ acts by holomorphic transformations. Such a holomorphic partial quotient is the quotient of the open subset $U_{\mathbb{C}}(\mathcal{K})$ by a torus-exponential action of the product of an algebraic torus and a complex space. In addition to complex moment-angle manifolds, this construction includes LVMB manifolds and complete nonsingular toric varieties as special cases.
@@ 12.1. Torus-exponential actions and their quotients
Given a vector configuration $\Gamma=\{\gamma_{1},\ldots,\gamma_{m}\}$ in $V$, the map $\varGamma^{*}|_{L}\otimes_{\mathbb{Z}}\mathbb{C}^{\times}$ embeds $\mathbb{C}^{\times}_{L}$ as a closed subgroup in (×$(\mathbb{C}^{\times})^{m}$:
@@
An action of a Lie group is almost free if all stabilisers are discrete subgroups.
@@
Let $\Gamma$ and $\mathrm{A}$ be a pair of Gale dual vector configurations, $Q\subset V$ a rational subspace, and $\mathcal{K}$ a simplicial complex on $[m]$. The torus-exponential action $(\mathbb{C}^{\times}_{L}\times_{Q}V)\times U_{\mathbb{C}}(\mathcal{K})\to U_{\mathbb{C}}(\mathcal{K})$ is proper if and only if $\{\mathcal{K},\mathrm{A}\}$ is a triangulated configuration, in which case the action is almost free with finite stabilisers.
@@
Since $T_{Q}$ is a compact group, the action of $\mathbb{C}^{\times}_{L}\times_{Q}V\cong T_{Q}\times V$ is proper if and only if the action of $V$ is proper. Hence, the assertion about the properness follows from Theorem 5.1. It also gives that the action of $V$ is free. For the action of $T_{L}$, the stabiliser of ${\textbf{{z}}}\in U_{\mathbb{C}}(\mathcal{K})$ is $2\pi(L_{I}/L)$. Since $I\in\mathcal{K}$, the set $\{\gamma_{k}\colon k\notin I\}$ spans $V^{*}$, so $L_{I}$ is a full-rank lattice in $Q$. It follows that $L_{I}/L$ is a finite group. ∎
@@
Let $\Gamma$ and $\mathrm{A}$ be a pair of Gale dual vector configurations, $Q\subset V$ a rational subspace, and $\mathcal{K}$ a simplicial complex on $[m]$. If $\{\mathcal{K},\mathrm{A}\}$ is a triangulated configuration, then the action of $\mathbb{C}^{\times}_{L}$ on $U_{\mathbb{C}}(\mathcal{K})$ is proper.
@@
is proper. This is the exponential action defined by $\widehat{\Gamma}=\{\gamma_{1},\ldots,\gamma_{m}\}$ viewed as a vector configuration in the quotient space $Q^{*}$ of $V^{*}$. Let $\widehat{\mathrm{A}}$ be the Gale dual configuration. Then $\{\mathcal{K},\widehat{\mathrm{A}}\}$ is a triangulated configuration by Proposition 11.3. It follows that the action is proper, by the criterion of Theorem 5.1. ∎
The quotient $U_{\mathbb{C}}(\mathcal{K})/\mathbb{C}^{\times}_{L}$ by a proper algebraic action of $\mathbb{C}^{\times}_{L}$ is the toric variety $X_{\widehat{\Sigma}}$ corresponding to the rational simplicial fan $\widehat{\Sigma}$ defined by $\{\mathcal{K},\widehat{\mathrm{A}}\}$ [16, §5.1].
@@ 12.2. Holomorphic torus-exponential actions
Let $\Gamma$ and $\mathrm{A}$ be a pair of Gale dual vector configurations, $Q\subset V$ a rational subspace, $\widehat{\Gamma}$ and $\widehat{\mathrm{A}}$ the corresponding rational configurations in $Q^{*}$ and $U^{*}$, and $\mathcal{K}$ a simplicial complex on $[m]$. Suppose that
@@
    $\{\mathcal{K},\mathrm{A}\}$ is a triangulated configuration defining a simplicial fan $\Sigma$;
@@
    $\{\mathcal{K},\widehat{\mathrm{A}}\}$ is a nonsingular triangulated configuration defining a nonsingular simplicial fan $\widehat{\Sigma}$ in $U^{*}$;
@@
Condition (a) implies that the action of $\mathbb{C}^{\times}_{L}\times\widetilde{V/Q}$ on $U_{\mathbb{C}}(\mathcal{K})$ is proper. Condition (b) together with Propositions 11.4 and 12.1 implies that the action of $T_{L}$ on $U_{\mathbb{C}}(\mathcal{K})$ is free, hence, the action of $\mathbb{C}^{\times}_{L}\times\widetilde{V/Q}$ is also free. Finally, the homeomorphism in the bottom right corner is proved in the same way as Theorem 10.6 (which corresponds to the case $Q=\{\mathbf{0}\}$). ∎
@@
If the rational triangulated configuration $\{\mathcal{K},\widehat{\mathrm{A}}\}$ is nonsingular, then the actions of $\mathbb{C}^{\times}_{L}$ and $\widetilde{T}_{L}$ are free, and $\mathcal{Z}_{\mathcal{K}}\to\mathcal{Z}_{\mathcal{K}}/\widetilde{T}_{L}$ is a holomorphic torus principal bundle over the holomorphic partial quotient. More generally, $\mathcal{Z}_{\mathcal{K}}\to\mathcal{Z}_{\mathcal{K}}/\widetilde{T}_{L}$ is a holomorphic Seifert principal bundle [37, §5.1] over the orbifold $\mathcal{Z}_{\mathcal{K}}/\widetilde{T}_{L}\cong X_{\widehat{\Sigma}}/(\widetilde{V/Q})$, where the latter the quotient of the toric orbifold $X_{\widehat{\Sigma}}$ by a free proper action of $\widetilde{V/Q}$.
@@ 12.3. Maximal torus actions
An action of a torus $T$ on a connected smooth manifold $M$ is maximal if there is a point $x\in M$ such that $\dim T+\dim T_{x}=\dim M$, where $T_{x}$ is the stabiliser subgroup of $x$. Examples of maximal torus actions are the action of a half-dimensional compact torus on a smooth complete toric variety, the free action of $T^{m}$ on itself and the action of $T^{m}$ on a moment-angle manifold $\mathcal{Z}_{\mathcal{K}}$. These are particular cases of the action of $T^{m}/T_{L}$ on a partial quotient $\mathcal{Z}_{\mathcal{K}}/T_{L}$:
@@
Compact complex manifolds $M$ with maximal torus actions by holomorphic transformations are classified by Ishida in [26]. The main result [26, Theorem 1.5] implies that every such $M$ is equivariantly biholomorphic to a holomorphic partial quotient $\mathcal{Z}_{\mathcal{K}}/T_{L}$. (For the latter formulation, see [44, Theorem 4.8].) It also implies that any complex manifold $M$ with a maximal torus action is the base of a holomorphic torus principal bundle with total space a complex moment-angle manifold.